Let f:[ 0,π/2] → [-1,1] be a function defined by f(x)= cos 2x . Verify that f satisfies the condition of the inverse function theorem. Hence, what can you conclude about the continuity of f^-1?

1
Expert's answer
2022-06-07T13:53:27-0400

The function "f(x)=\cos2x" is strictly decreasing on "(0, \pi/2)." Then the function "f(x)=\cos2x" is one-to-one on "[0, \pi/2]."

The function "f(x)=\cos2x" is invertible on "[0, \pi/2]."

"f'(x)=-2\sin(2x)"

By  Inverse Function Theorem for all "x\in(0, \pi/2)"


"(f^{-1}(x))'=\dfrac{1}{f'(f^{-1}(x))}"

"f^{-1}(x)" is continuous on "[-1,1]."


"f(x)=\cos2x, x\in [0, \pi/2]"

"f^{-1}(x)=\dfrac{1}{2}\cos^{-1}x, x\in[-1,1]"


"(f^{-1}(x))'=(\dfrac{1}{2}\cos^{-1}x)'=-\dfrac{1}{2\sqrt{1-x^2}}"


"\dfrac{1}{f'(f^{-1}(x))}=\dfrac{1}{-2\sin(2(\dfrac{1}{2}\cos^{-1}x))}=-\dfrac{1}{2\sqrt{1-x^2}}"